OpenStudy (anonymous):

what steps would I take to explain these the difference in these 3 problems a2 - 25 a2 + 8a + 16 a2 - 8a + 16

OpenStudy (annon):

is that a^2 or 2a?

OpenStudy (anonymous):

a^2-25 =(a)^2-(5)^2 (a+5)(a-5)

OpenStudy (anonymous):

first one is not aperfect square . and the remaining are perfect squares

OpenStudy (anonymous):

a^2+2+4a+(4)^2 =(a+4)^2

OpenStudy (anonymous):

a^2-2*4a+(4)^2 =(a-4)^2

OpenStudy (anonymous):

third one (a-4)^2

OpenStudy (anonymous):

it's a to the second power